# THOR2: Leveraging Topological Soft Clustering of Color Space for Human-Inspired Object Recognition in Unseen Environments

### v1

@@ 6.3.1 Instance Segmentation Quality
The TOPS and TOPS2 descriptors embody the idea of object unity, making THOR2 relatively robust to over-segmentation errors or errors where the object masks are split into segments due to occlusion. However, under-segmentation and false positive segmentations are more challenging. Further, a certain degree of robustness to errors arising from misaligned object boundaries is also achieved through the test-time outlier removal from point clouds. On a related note, the temporally smoothed depth images (and point clouds) provided in the OCID dataset are relatively less noisy and do not require outlier removal. Such temporal smoothing leads to a misalignment between the RGB and depth image corresponding to a given timestep in the sequence. In sequences where the misalignment is more pronounced (e.g., sequences S-02 and S-07), the object point clouds are incorrectly colored on the boundaries resulting in a somewhat lower performance using the shape and color-based THOR2 compared to exclusively shaped-based THOR.
@@ 7 Conclusion
We report performance evaluation on two benchmark datasets: OCID and UW-IS Occluded. On the OCID dataset, comparisons with THOR show that THOR2 benefits from incorporating color information. Moreover, THOR2 achieves the best overall performance among all the deep learning-based end-to-end models, including the widely-used transformer, ViT, adapted for RGB-D input when increasingly cluttered scenes are viewed from different camera positions. THOR2 also achieves substantially higher recognition accuracy on the UW-IS Occluded dataset than all the methods, regardless of the environmental condition. Moreover, THOR2 continues to outperform RGB-D ViT even when real-world RGB-D images from the YCB dataset are used along with synthetic images to train the latter.

### v2

@@ 6.4.1 Instance Segmentation Quality
The TOPS and TOPS2 descriptors embody the idea of object unity, making THOR2 relatively robust to over-segmentation errors or errors where the object masks are split into segments due to occlusion. However, under-segmentation and false positive segmentations are more challenging. Further, a certain degree of robustness to errors arising from misaligned object boundaries is also achieved through the test-time outlier removal from point clouds. On a related note, the temporally smoothed depth images (and point clouds) provided in the OCID dataset are relatively less noisy and do not require outlier removal. Such temporal smoothing leads to a misalignment between the RGB and depth image corresponding to a given timestep in the sequence. In sequences where the misalignment is more pronounced (e.g., sequences S-02 and S-07), the object point clouds are incorrectly colored on the boundaries resulting in a somewhat lower performance using the shape and color-based THOR2 as compared to exclusively shaped-based THOR.
@@ 7 Conclusion
We report performance evaluation on two benchmark datasets: OCID and UW-IS Occluded. On the OCID dataset, comparisons with THOR show that THOR2 benefits from incorporating color information. Moreover, THOR2 achieves the best overall performance among all the deep learning-based end-to-end models, including the widely-used transformer, ViT, adapted for RGB-D input when increasingly cluttered scenes are viewed from different camera positions. THOR2 also achieves substantially higher recognition accuracy on the UW-IS Occluded dataset than all the methods, regardless of the environmental condition. Moreover, THOR2 continues to outperform deep learning-based end-to-end models even when real-world RGB-D images from the YCB dataset are used along with synthetic images to train the models.